# Impact of Annealing and Nanostructuring on Properties of NV Centers Created by Different Techniques

### v1

@@ III.5 Electron Spin Coherence
To measure the coherence times, a Hahn-echo sequence hahn was applied to the electronic spin of the NVs, using the lowest ODMR resonant frequency from 2.77 GHz to 2.78 GHz. Rabi measurements rabi were performed prior to the Hahn-echo to obtain the exact $\pi$-pulse duration. Finally, the $T_{2}$ times were obtained by fitting the Hahn-echo decay with an exponential function, with an error corresponding to the standard deviation from the fit parameter. The results for the bulk regions before and after annealing are shown in Fig. 7.
Annealing does not influence the coherence in the (111)-implanted sample, while the $\delta$-doped shows a small increase, but still within the uncertainty. Only the (100)-implanted sample has a definite increase of $T_{2}$ by a factor of 1.4. Regarding the surface orientation, the (111)-implanted sample has much shorter coherence than the (100) sample. Furthermore, the latter one and the $\delta$-doped samples have similar $T_{2}$ values. The coherence times in the pillars before annealing could not be measured due to insufficient optical contrast between the $m_{s}$ states in the $\delta$-doped sample and bleaching effects in the implanted ones. After annealing (data not shown), $T_{2}$ are comparable to the bulk with typically larger uncertainty, which could be related to magnetic noise coming from surface spins in the pillar walls.
@@ IV Conclusions
This work contributes to the ongoing advancement and characterization of the fabrication techniques of color centers in diamond and potentially other systems in quantum technologies.
Further studies can concentrate on the very high fluorescence enhancements in the smallest pillars or the determination of the exact concentration of NV and P1 centers in differently prepared samples by double electron-electron resonance (DEER) techniques DEER .

### v2

@@ III.5 Electron Spin Coherence
To measure the coherence times, a Hahn-echo sequence Hahn (1950) was applied to the electronic spin of the NVs, using the lowest ODMR resonant frequency from 2.77 GHz to 2.78 GHz. Rabi measurements Rabi et al. (1939) were performed prior to the Hahn-echo to obtain the exact $\pi$-pulse duration. Finally, the $T_{2}$ times were obtained by fitting the Hahn-echo decay with an exponential function, with an error corresponding to the standard deviation from the fit parameter. The results for the bulk regions before and after annealing are shown in Fig. 7.
The (111)-implanted sample has much shorter coherence than the (100)-implanted and $\delta$-doped samples which show similar $T_{2}$ values. A possible explanation could be the presence of some electron spin bath (for example P1 centers with electron spin $S=1/2$) in the diamond crystal even before the implantation which cannot be removed during the annealing step. The latter induced a small increase of $T_{2}$ in the $\delta$-doped sample (still within the uncertainty) and a definite increase by a factor of 1.4 in the (100)-implanted one.
The coherence times in the pillars before annealing could not be measured due to insufficient optical contrast between the $m_{s}$ states in the $\delta$-doped sample and bleaching effects in the implanted ones. After annealing (data not shown), $T_{2}$ are comparable to the bulk with typically larger uncertainty, which could be related to magnetic noise coming from surface spins in the pillar walls.
@@ IV Conclusions
This work contributes to the ongoing advancement and characterization of the fabrication techniques of color centers in diamond and potentially other systems in quantum technologies. Specifically, these results constitute an initial step towards the preparation of recently proposed diamond based quantum token Tsunaki et al. (2024b); Bauerhenne et al. (2024), aiming to use quantum states of NV ensembles in nanopillars as unclonable authentication keys. Further studies can concentrate on the very high fluorescence enhancements in the smallest pillars or the determination of the exact concentration of NV and P1 centers in differently prepared samples by double electron-electron resonance (DEER) techniques Li et al. (2021).
@@ Acknowledgments
Sample A [nm] B [nm] C [nm] D [nm] E [nm]
nominal diameters 300 250 200 150 100
$\delta$-dopped $280\pm 30$ $240\pm 20$ $190\pm 10$ $160\pm 10$ $100\pm 10$
(100)-implanted $280\pm 20$ $250\pm 10$ $200\pm 10$ $150\pm 20$ $142\pm 9$
(111)-implanted $290\pm 20$ $240\pm 10$ $190\pm 10$ $142\pm 8$ $130\pm 20$
Table S1: Average diameters of the nanopillars determined from the SEM images for each sample. Uncertainty is given by the standard deviation.
@@
Figure S3: Area normalized saturation curves of $\delta$-doped sample measured in the pillars before and after each annealing step. The area normalized fluorescence shows a great increase in the smallest pillars, as also shown in Fig. 5 (d).
Figure S4: Area normalized saturation curves of the implanted sample measured in the pillars after annealing at 1400°C. The (111)-implanted has a small increase in fluorescence in the smaller pillars, while the (100)-shows a decrease.